OpenStudy (anonymous):

solve for t ln 4.4=ln e Power (.031t)

OpenStudy (anonymous):

is this \[\ln(4.4)=\ln(e^{.031t})\]

OpenStudy (anonymous):

right hand side is \[.031t\] since \[\ln(e^x)=x\]

OpenStudy (anonymous):

so you really just have \[\ln(4.4)=.031t\] \[t=\frac{\ln(4.4)}{.031}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

does my calculator need to be in a certain mode to use the ln function?